# End-to-end Learned Visual Odometry with Events and Frames

### v2

@@ Abstract
Visual Odometry (VO) is crucial for autonomous robotic navigation, especially in GPS-denied environments like planetary terrains. To improve robustness, recent model-based VO systems have begun combining standard and event-based cameras. Event cameras excel in low-light and high-speed motion, while standard cameras provide dense and easier-to-track features, even in low-textured areas. However, the field of image- and event-based VO still predominantly relies on model-based methods and is yet to fully integrate recent image-only advancements leveraging end-to-end learning-based architectures. Seamlessly integrating the two modalities remains challenging due to their different nature, one asynchronous, the other not, limiting the potential for a more effective image- and event-based VO. We introduce RAMP-VO, the first end-to-end learned image- and event-based VO system. It leverages novel Recurrent, Asynchronous, and Massively Parallel (RAMP) encoders capable of fusing asynchronous events with image data, providing 8×8$\times$8 × faster inference and $33\%$ more accurate predictions than existing solutions. Despite being trained only in simulation, RAMP-VO outperforms image- and event-based methods by $46\%$ and $60\%$, respectively, on traditional, real-world benchmarks as well as newly introduced Apollo and Malapert landing sequences, paving the way for robust and asynchronous VO in space.
@@ I Introduction
Visual Odometry (VO) is a central building block in many robotic platforms. Traditionally, VO systems have relied exclusively on standard RGB cameras to estimate the pose accurately. However, when navigating in challenging scenarios, such as low-light environments, high dynamic range scenes, or low-textured terrains at high speed, standard VO methods typically fail. These shortcomings are mostly caused by their susceptibility to motion blur, the limited dynamic range, as well as their unfavorable bandwidth-vs-latency tradeoff. While a higher frame rate would reduce latency, it comes at the cost of a higher bandwidth and increased processing power. Event-based cameras, on the other hand, promise to address all these issues. They are bio-inspired sensors that asynchronously record per-pixel brightness changes at microsecond resolution. They exhibit high dynamic range (HDR), low latency, and low power consumption, making them an ideal complement to regular cameras in VO systems.
Their combination holds significant promise for critical VO applications, especially in scenarios where conventional sensors like GPS, LiDAR, and Inertial Measurement Units (IMUs) cannot be used or are ineffective due to radiation and temperature changes. These conditions are often encountered in planetary exploration and landing missions, where rapid motion and partial shadow are also common.
@@
Figure 1: Overview of the proposed method: Recurrent, Asynchronous, and Massively Parallel (RAMP) Encoders are used to process asynchronous events and images. Patches are extracted from the resulting encoding and used by the Estimator inspired by DPVO [1] to perform data-driven feature tracking and visual odometry. A simple pose forecasting module exploits previously extracted patches to initialize poses in the bundle adjustment, allowing for improved performance.
@@
Recent model-based solutions [2] have shown potential in fusing images and events. However, the fields of image-only [1, 3] and event-only [4] VO have recently shown that learning-based pipelines, trained end-to-end, have the potential to surpass traditional model-based systems in accuracy and robustness. While the combination of data-driven approaches with systems that leverage images and events appears promising, effectively combining event data—with its distinctive asynchronicity and sparsity—with synchronous and dense frames is a non-trivial challenge in learning-based solutions. Traditional learning-based methods typically resort to artificially synchronizing events at image timestamps to facilitate data fusion, reducing the rate at which events are processed to that of the slower image modality. Nevertheless, in tasks such as VO, this simplification is not ideal and might limit the algorithm’s ability to exploit events received in between images, which is crucial for tracking features effectively.
@@
To address these limitations, our work introduces an adaptive fusion approach that adjusts the frequency of event fusion based on the rate of incoming events, thus mirroring the pace of the scene dynamics. Our Recurrent, Asynchronous, and Massively Parallel (RAMP) encoders handle asynchronous [^1] Notice that, in this work, with the term “asynchronous” we refer to the ability of our network to handle data streams operating at different, and possibly varying, rates. Our network processes either images or packets of events (in the form of frame-like event representations) as soon as they are available, without a stream synchronization step. This should not be confused with networks operating on event-by-event processing where “asynchronous” refers to the network’s layers functioning. events and images at varying rates and fuse them into a pyramidal memory that serves as a data-agnostic feature space. We use these encoders in RAMP-VO, the first learning-based VO method that uses both events and frames. RAMP-VO leverages a motion-aware strategy based on event data to extract robust patch-based feature tracks. These features are then processed by a differentiable bundle adjustment module [1] which leverages a simple pose forecasting module for initialization.
@@
  • A novel massively parallel feature extractor, termed RAMP encoder that fuses images and events, both spatially and temporally. Our encoder is 8 times faster and achieves a 20% higher performance than other state-of-the-art asynchronous solutions.
  • RAMP-VO, the first learning-based VO using events and frames, which outperforms both image-based and event-based methods by $46\%$ and $60\%$, respectively on traditional real-world benchmarks.
@@ III Methodology
Figure 2: An overview of the proposed RAMP Net encoder. Events and images are first asynchronously processed by two parallel pixel-wise, multi-scale, encoding branches (PWE) made of a set of convolutional layers followed by pixel-wise LSTMs $G_{k}^{s}$. Features coming from different data modalities $k$ are then fused (SF), at each scale, into a shared state $\Sigma^{s}_{t}$ by employing sensor-specific encoders. The multi-scale features are then finally combined through two separate fusion modules (MSF) to produce the matching and context features, $m_{t}$ and $c_{t}$.
@@ III-A Overview
Our end-to-end event- and frame-based visual odometry algorithm, RAMP-VO, takes inspiration from recurrent asynchronous multimodal (RAM) networks [44] and deep patch visual odometry (DPVO) [1]. Our RAMP encoder processes a temporally ordered, asynchronous stream of data, which we will call frames. These are either standard images or frames made of events. Similar to DPVO [1], for each new frame with index $j$ we first compute matching $m_{j}$ and context features $c_{j}$ through RAMP encoder, described in III-B, and then extract $N$ patches with dimension $p\times p$ from these feature maps.
While DPVO [1] and DEVO [4] select these patches randomly or through a learning-based strategy, respectively, we opt for a simpler yet effective alternative based on events’ activity. In particular, we first compute an $H\times W$ density map by counting the number of events for each pixel. Then, we select the $N$ locations on this map with the highest counts, after having applied non-maximum suppression, which removes candidates that are not the highest in their $11\times 1$ 11 local neighborhood. The thus sampled locations act as the centers of the selected patches.
Denote a patch $\mathbf{P}^{l}_{j}$ with index $l$ extracted in frame $j$ as
$\mathbf{P}^{l}_{j}=\begin{bmatrix}\mathbf{x}&\mathbf{y}&\mathbf{1}&\mathbf{d}\end{bmatrix}^{T}\quad\mathbf{x},\mathbf{y},\mathbf{d}\in\mathbb{R}^{1\times p^{2}}.$ (1)
@@
These patches are then projected into the previous frames of index $i\in\{j-r,j-r+1,...,j\}$ and patches extracted in previous frames (going back to frame $j-r$) are projected into the current frame $j$. A visualization is provided in Figure 3. Let the projection of patch $P^{l}_{j}$, into frame $i$ be
$\mathbf{P^{\prime}}^{l}_{ji}\sim\mathbf{K}\mathbf{T}_{i}\mathbf{T}_{j}^{-1}\mathbf{K}^{-1}\mathbf{P}^{l}_{j}$ (2)
where $K$ is the $4\times 4$ camera matrix and $T_{i},T_{j}$ are poses at frames $i,j$. We summarize this as $\mathbf{P^{\prime}}^{l}_{ji}=\omega(\mathbf{T}_{i},\mathbf{T}_{j},\mathbf{P}^{l}_{j})$.
Next, RAMP-VO computes camera motion by estimating 2D corrections $\Delta_{li}\in\mathbb{R}^{2}$ for each projected patch $\mathbf{P^{\prime}}^{l}_{ji}$, as well as importance values $\sigma_{li}\in\mathbb{R}^{2\times 2}$ through a series of blocks involving a correlation lookup, 1D Convolution, Soft Aggregation, Transition Block and Factor Head. Since these operations are out of the scope of the current work, we summarize them with the following relation
$\Delta_{li},\sigma_{li}=F(\mathbf{P^{\prime}}^{l}_{ji},c_{i},m_{i},m_{j})$ (3)
where we compare context features $c_{i}$ with matching features $m$ extracted in frame $i$ and $j$. The operator $F$ represents the update operator from [1], represented inside the estimator block in Fig. 1.
@@
Here the left-hand side is kept fixed, while the optimization solves for the camera poses $\mathbf{T}_{i},\mathbf{T}_{j}$ and depths $\mathbf{d}^{l}_{j}$. The $\hat{}$ operator selects the central pixel of the patch, while $\mathcal{E}$ is a bipartite graph connecting frames with patches projected onto each of them. By using two Levenberg-Marquart optimization steps, the BA layer updates the depth and camera poses in the time window. This operation is fully differentiable, and thus used during training to backpropagate errors.
Contrary to [1] which uses a simple linear initialization, we bootstrap initial poses in our RAMP-VO by fitting a continuous-time model over the available patch tracks, as depicted in Figure 3. We do so by fitting two cubic univariate splines $S^{l}_{x}(t)$ and $S^{l}_{y}(t)$ modeling the 2D motion of the patch center along the last $11$ frames, such that $(\mathbf{x}^{l}_{i},\mathbf{y}^{l}_{i})\sim\mathbf{S}^{l}(t_{i})=(S^{l}_{x}(t_{i}),S^{l}_{y}(t_{i}))$. We then extrapolate the location of the camera at time $t_{j+1}$, by first evaluating $\mathbf{S}^{l}(t_{j+1})$, assuming the first and second derivatives to be zero at the splines’ boundaries and the depth to remain constant, and then optimizing its 6 DOF pose through BA:
$\sum_{(l,i)\in\mathcal{E}}\left\|\mathbf{S}^{l}(t_{j+1})-\hat{\omega}(T_{i},T_{j},\mathbf{P}^{l}_{j})\right\|^{2}_{\sigma_{li}}.$ (5)
The main innovations of this work are related to how the context and matching features $c_{j}$ and $m_{j}$ are generated by fusing asynchronous events and frames. These topics are discussed next.
@@ III-B Asynchronous and Massively Parallel Encoders
Denote the stream of data $\{x_{k_{j}}(t_{j})\}_{j=1}^{T}$ captured at timestamps $t_{j}$. Here $x$ (henceforth denoted as ”frame”) denotes either a $5\times H\times W$ sized event stack  [45], in the case of events, or a $C\times H\times W$ sized image [^2] For color images $C=3$ and for gray-scale images $C=1$.. The variable $k_{j}\in\{\text{e},\text{i}\}$ denotes the sensor at timestamp $t_{j}$. We encode these data structures using a Recurrent, Asynchronous and Massively Parallel (RAMP) encoder. An overview of the architecture is provided in Figure 2. We employ two different Multi-Scale Fusion modules, with identical structures, to generate either the context or the matching feature maps.
The RAMP encoders first transduce the data stream using pixel-wise encoders starting with a stream of features $\{f^{s}_{k_{j}}\}_{j=1}^{T}$ at three different scales. To encode this data we use separate sensor-specific encoders, one for each scale
$f^{s}_{k_{j}}(t_{j})=E^{s}_{k_{j}}(x_{k_{j}}(t_{j})).$ (6)
@@
This stream of features is then further encoded in two recurrent stages: (1) intra-sensor fusion, and (2) inter-sensor fusion. In the intra-sensor fusion step, we process features originating from a single sensor and scale with an LSTM operating on individual pixels, inspired by [46]
$h^{s}_{k_{j}}(t_{j})=G^{s}_{k_{j}}(f^{s}_{k_{j}}(t_{j})),$ (7)
where we have omitted the cell state for the sake of clarity.
The inter-sensor fusion step, instead, fuses the hidden states from separate sensors asynchronously using only a single depth-wise convolution in the following way:
$\Sigma^{s}_{j}=H_{k_{j}}\left([h^{s}_{k_{j}}(t_{j})\|\Sigma_{j-1}]\right)$ (8)
@@
Note that the RAMP encoder fuses each pixel independently and only relies on a simple inter-sensor fusion strategy, and thus can be done efficiently in parallel. This is in contrast to the encoder in  [47], which requires sequential processing, and expensive ConvGRU operations at two stages. After generating inter-sensor fused states $\Sigma_{j}$ at each time $t_{j}$, a hierarchical encoder is used to generate either the context or the matching feature maps, each at $1/4$ the original resolution
$z_{j}=K(\{\Sigma_{j}^{s}\}_{s=0}^{2}).$ (9)
Here $z_{j}$ can be either $c_{j}$ or $m_{j}$, i.e., context and matching features, required by the DPVO backend.
Multi-Scale Fusion.   The Multi-scale Fusion (MSF) module follows the feature extractors in [1] and consists of a $7\times 7$ convolution with stride $2$, four residual blocks (two at $1/2$ and two at $1/4$ of the initial resolution), and a final $1\times 1$ convolution. We use the same configuration in [1], providing full-scale features $\Sigma_{j}^{0}$ as input to the encoder, but injecting $\Sigma_{j}^{1}$ and $\Sigma_{j}^{2}$ after the second and fourth residual blocks respectively. We do so by concatenating these features with the residual block’s outputs and adjusting the channels of the next operation to accommodate the additional features.

### v3

# Deep Visual Odometry with Events and Frames

@@ Abstract
Visual Odometry (VO) is crucial for autonomous robotic navigation, especially in GPS-denied environments like planetary terrains. To improve robustness, recent model-based VO systems have begun combining standard and event-based cameras. While event cameras excel in low-light and high-speed motion, standard cameras provide dense and easier-to-track features. However, the field of image- and event-based VO still predominantly relies on model-based methods and is yet to fully integrate recent image-only advancements leveraging end-to-end learning-based architectures. Seamlessly integrating the two modalities remains challenging due to their different nature, one asynchronous, the other not, limiting the potential for a more effective image- and event-based VO. We introduce RAMP-VO, the first end-to-end learned image- and event-based VO system. It leverages novel Recurrent, Asynchronous, and Massively Parallel (RAMP) encoders capable of fusing asynchronous events with image data, providing 8×8$\times$8 × faster inference and $33\%$ more accurate predictions than existing solutions. Despite being trained only in simulation, RAMP-VO outperforms previous methods on the newly introduced Apollo and Malapert datasets, and on existing benchmarks, where it improves image- and event-based methods by $58.8\%$ and $30.6\%$, paving the way for robust and asynchronous VO in space.
Multimedial Material: For code and datasets visit https://github.com/uzh-rpg/rampvo.
@@ I Introduction
Visual Odometry (VO) is vital for robotic platforms but often fails in challenging scenarios involving low-light, high dynamic range, or high-speed motion. These shortcomings are often caused by the limitations of traditional frame-based camera, such as their susceptibility to motion blur, limited dynamic range, and unfavorable bandwidth-vs-latency tradeoff. While higher frame rates reduce latency, they come at the cost of higher bandwidth and increased processing power. Event-based cameras, which record per-pixel brightness changes asynchronously, address all these issues, offering high dynamic range (HDR), low latency, and low bandwidth and power usage, making them the ideal complement to regular cameras in VO systems. Their combination holds significant promise for critical VO applications, especially when sensors like GPS, LiDAR, and Inertial Measurement Units (IMUs) cannot be used or are ineffective due to radiation and temperature changes. These conditions are often encountered in planetary exploration and landing missions, where rapid motion and partial shadow are also common.
@@
Figure 1: Recurrent, Asynchronous, and Massively Parallel (RAMP) Encoders are used to process asynchronous events and images. Patches are extracted from the resulting encoding and used by the Estimator inspired by DPVO [1] to perform data-driven feature tracking and visual odometry. A simple pose forecasting module exploits previously extracted patches to initialize poses in the bundle adjustment, allowing for improved performance.
@@
Recent model-based solutions [2] have shown potential in fusing images and events. However, the fields of image-only [1, 3] and event-only [4] VO have recently shown that learning-based pipelines, trained end-to-end, can surpass traditional model-based systems in accuracy and robustness. While the combination of data-driven approaches with systems that leverage images and events appears promising, effectively combining event data—with its distinctive asynchronicity and sparsity—with synchronous and dense frames is a non-trivial challenge in learning-based solutions. Traditional learning-based methods typically resort to artificially synchronizing events at image timestamps to facilitate data fusion, reducing the rate at which events are processed to that of the slower image modality. Nevertheless, in tasks such as VO, this simplification is not ideal and might limit the algorithm’s ability to exploit events received in between images, which is crucial for tracking features effectively.
@@
To address these limitations, our work introduces an adaptive fusion approach that adjusts the frequency of event fusion based on the rate of incoming events, thus mirroring the pace of the scene dynamics. Our Recurrent, Asynchronous, and Massively Parallel (RAMP) encoders handle asynchronous events and images at varying rates and fuse them into a pyramidal memory that serves as a data-agnostic feature space. We use these encoders in RAMP-VO, the first learning-based VO method that uses both events and frames. RAMP-VO leverages a motion-aware strategy based on event data to extract robust patch-based feature tracks. These features are then processed by a differentiable bundle adjustment module [1] which leverages a simple pose forecasting module for initialization.
@@
  • A novel massively parallel feature extractor, termed RAMP encoder, that asynchronously [^1] Notice that in this work, the term “asynchronous” refers to our network’s ability to handle data streams at different and varying rates. Our network processes images or packets of events (as frame-like event representations) as soon as they are available, without stream synchronization. This differs from networks operating on event-by-event processing, where “asynchronous” refers to the network layers’ functioning. fuses images and events, both spatially and temporally. Our encoder is 8 times faster and achieves a 33% higher performance than other state-of-the-art asynchronous solutions.
  • RAMP-VO, the first learning-based VO using events and frames, which outperforms both image-based and event-based methods by $58.8\%$ and $30.6\%$, respectively on traditional real-world benchmarks.
@@ III Methodology
Figure 2: An overview of the proposed RAMP Net encoder. Events and images are first asynchronously processed by two parallel pixel-wise, multi-scale, encoding branches (PWE) made of a set of convolutional layers followed by pixel-wise LSTMs $G_{k}^{s}$. A shared state $\Sigma^{s}_{t}$ is then updated (SU) with features coming from different data modalities $k$ by employing sensor-specific encoders at each scale. The multi-scale features are then finally combined through two separate fusion modules (MSF) to produce the matching and context features, $m_{t}$ and $c_{t}$
Our end-to-end event- and frame-based visual odometry algorithm, RAMP-VO, builds on deep patch visual odometry (DPVO) [1] and takes inspiration from recurrent asynchronous multimodal (RAM) networks [44]. The main innovations of this work reside in how events and frames are fused together. An ideal vision encoder should fuse sensor measurements at an adaptable rate, processing more information during fast motion for better tracking, and ensuring reliability even when a sensor experiences temporary outages. Our RAMP encoders achieve this by creating a sensor-agnostic representation that is recurrently updated whenever new data becomes available, either images or events. We present the main components of RAMP-VO in Section III-A, and the RAMP encoder in Section III-B.
@@ III-A Overview
RAMP-VO processes a temporally ordered, asynchronous stream of data made of ‘frames’. These are either standard images or frames made of events with $H\times W$ resolution. Similar to DPVO [1], we extract a set of patches and predict their optical flow across frames by computing correlation volumes over matching features $\mathbf{m}_{j}$ and additional context features $\mathbf{c}_{j}$. We iteratively refine the predicted patch flow and finally use a differentiable bundle adjustment layer to update the camera poses. For each new frame with index $j$ we first compute $\mathbf{m}_{j}$ and $\mathbf{c}_{j}$ through the RAMP encoder described in Section III-B, and then extract $N$ patches with dimension $p\times p$ from these feature maps. While DPVO [1] and DEVO [4] select these patches randomly or through a learning-based strategy, we opt for a simpler yet effective alternative. In particular, we rely on events’ activity to extract patches from textured areas, as these regions usually generate a higher number of events compared to uniform ones. We first compute an $H\times W$ density map by counting the number of events per pixel. We then apply non-maximum suppression to remove candidates that are not the highest in their $11\times 1$ 11 neighborhood and finally obtain the patch centers by selecting the $N$ locations with the highest counts.
We denote the $l$-th patch $\mathbf{P}^{l}_{j}$ extracted from frame $j$ as
$\mathbf{P}^{l}_{j}=\begin{bmatrix}\mathbf{x}&\mathbf{y}&\mathbf{1}&\mathbf{d}\end{bmatrix}^{T},\quad\mathbf{x},\mathbf{y},\mathbf{d}\in\mathbb{R}^{1\times p^{2}}.$ (1)
@@
Following [1], we build a bipartite patch graph $\mathcal{E}$, depicted in Figure 3, by connecting a patch $\mathbf{P}^{l}_{j}$ extracted from frame $j$ to every frame $i$ within a distance $r$ from $j$, each containing the projection $\mathbf{P^{\prime}}^{l}_{ji}$ of the original patch onto the frame:
$\mathbf{P^{\prime}}^{l}_{ji}\sim\mathbf{\bar{K}}\mathbf{T}_{i}\mathbf{T}_{j}^{-1}\mathbf{\bar{K}}^{-1}\mathbf{P}^{l}_{j},\quad\mathbf{\bar{K}}=\begin{pmatrix}\mathbf{K}&\mathbf{0}\\ \mathbf{0}^{T}&1\end{pmatrix}$ (2)
where $\mathbf{\bar{K}}$ is a $4\times 4$ matrix built from the $3\times 3$ camera matrix $\mathbf{K}$ and $\mathbf{T}_{i},\mathbf{T}_{j}$ are poses of frames $i,j$. We summarize this operation as $\mathbf{P^{\prime}}^{l}_{ji}=\omega(\mathbf{T}_{i},\mathbf{T}_{j},\mathbf{P}^{l}_{j})$.
Next, RAMP-VO computes camera motion by estimating 2D corrections $\boldsymbol{\Delta}_{li}\in\mathbb{R}^{2}$ for each projected patch $\mathbf{P^{\prime}}^{l}_{ji}$, as well as importance values $\sigma_{li}\in\mathbb{R}^{2\times 2}$ through a series of blocks involving a correlation lookup, 1D Convolution, Soft Aggregation, Transition Block and Factor Head. These steps extract the features of each patch in frame $j$ from $\mathbf{m}_{j}$ and compare them with those obtained by cropping $\mathbf{m}_{i}$ with the reprojection of the patch in frame $i$, together with additional context features $\mathbf{c}_{i}$. Since these operations, described in [1], are out of the scope of the current work, we summarize them with the update operator $F$ in Fig. 1:
$\Delta_{li},\sigma_{li}=F(\mathbf{P^{\prime}}^{l}_{ji},\mathbf{c}_{i},\mathbf{m}_{i},\mathbf{m}_{j})$ (3)
@@
Here the left-hand side is kept fixed, while the optimization solves for the camera poses $\mathbf{T}_{i},\mathbf{T}_{j}$ and depths $\mathbf{d}^{l}_{j}$, refining the camera trajectory to match the predicted patch trajectories. The $\hat{}$ operator selects the central pixel of the patch, and $\mathcal{E}$ is the path graph. This operation is fully differentiable, and thus used during training to backpropagate errors.
Contrary to [1] which uses a simple linear initialization, we bootstrap initial poses in our RAMP-VO by fitting two cubic univariate splines modeling the 2D motion of the patch center along the last $11$ frames, such that $(\mathbf{x}^{l}_{j},\mathbf{y}^{l}_{j})\sim\mathbf{S}^{l}(t_{j})=(S^{l}_{x}(t_{j}),S^{l}_{y}(t_{j}))$. We then extrapolate the location of the camera at time $t_{j+1}$, by evaluating $\mathbf{S}^{l}(t_{j+1})$, assuming the first and second derivatives to be zero at the splines’ boundaries and constant depth. We obtain the 6 DOF pose at time $t_{j+1}$ through BA, optimizing Equation 4 in which we use $\mathbf{S}^{l}(t_{j+1})$ in place of the corrected patch position.
@@ III-B Asynchronous and Massively Parallel Encoders
Denote the stream of data $\{\mathbf{I}_{k_{j}}(t_{j})\}_{j=1}^{T}$ captured at timestamps $t_{j}$. Here $\mathbf{I}$ (henceforth denoted as ”frame”) denotes either a $5\times H\times W$ sized event stack  [45], in the case of events, or a $C\times H\times W$ sized image [^2] For color images $C=3$ and for gray-scale images $C=1$.. The variable $k_{j}\in\{\text{e},\text{i}\}$ denotes the sensor at timestamp $t_{j}$. We encode these data structures using a Recurrent, Asynchronous and Massively Parallel (RAMP) encoder. An overview of the architecture is provided in Figure 2. We employ two different Multi-Scale Fusion modules, with identical structures, to generate either the context or the matching feature maps.
We designed the RAMP encoders to be easily parallelizable and to effectively fuse asynchronous data streams. To do so, we encode each pixel independently, in contrast with  [46] which requires sequential processing and expensive two-stage ConvGRU operations, and recurrently fuse information into a shared state. First, RAMP encoders process the data stream with pixel-wise, sensor-specific networks at multiple scales $s$, creating a sequence of features $\{\mathbf{f}^{s}_{k_{j}}\}_{j=1}^{T}$:
$\mathbf{f}^{s}_{k_{j}}(t_{j})=E^{s}_{k_{j}}(\mathbf{I}_{k_{j}}(t_{j})).$ (5)
@@
This stream of features is then further encoded into a sensor-agnostic state through two recurrent stages: (1) intra-sensor update, and (2) inter-sensor update. In the intra-sensor update step, we process features originating from a single sensor and scale using an LSTM operating on individual pixels, inspired by [47]
$\mathbf{h}^{s}_{k_{j}}(t_{j})=G^{s}_{k_{j}}(\mathbf{f}^{s}_{k_{j}}(t_{j})),$ (6)
where we have omitted the cell state for brevity.
The inter-sensor update step, finally combines together the hidden states from separate sensors asynchronously using only a single depth-wise convolution in the following way:
$\boldsymbol{\Sigma}^{s}_{j}=H_{k_{j}}\left([\mathbf{h}^{s}_{k_{j}}(t_{j})\|\Sigma_{j-1}]\right)$ (7)
@@
After generating the sensor-agnostic states $\Sigma_{j}$ at each time $t_{j}$, two hierarchical encoders generate either the context or the matching feature maps, each at $1/4$ the original resolution
$\mathbf{z}_{j}=K(\{\Sigma_{j}^{s}\}_{s=0}^{2}).$ (8)
Here $\mathbf{z}_{j}$ can be either $\mathbf{c}_{j}$ or $\mathbf{m}_{j}$, i.e., the context and matching features required by the DPVO backend.
Multi-Scale Fusion.   The Multi-scale Fusion (MSF) module follows the single-scale feature extractors in [1] and consists of a $7\times 7$ convolution with stride $2$, four residual blocks (two at $1/2$ and two at $1/4$ of the initial resolution), and a final $1\times 1$ convolution. We follow the configuration in [1], using full-scale features $\Sigma_{j}^{0}$ as input to the encoder, but additionally injecting $\Sigma_{j}^{1}$ and $\Sigma_{j}^{2}$ after the second and fourth residual blocks respectively. We do so by concatenating these features to the residual block’s outputs and adjusting the channels of the next operation to accommodate the additional features.